# Unlearning Inversion Attacks for Graph Neural Networks

### v1

@@ D.3 Edge Influence on Model Output
Theorem D.9 (Edge Influence on Model Output, Linear GCN Case).
Let $\boldsymbol{\Xi}^{\Delta\mathcal{E}}\in\mathbb{R}^{n\times n}$ be a perturbation matrix as defined in Definition D.6. Let $\epsilon\in\mathbb{R}$ be a perturbation magnitude and $\mathbf{w}^{\star}(\epsilon)$ be the optimal model weight after perturbation. Considering an edge perturbation $\mathbf{C}(\epsilon):=\mathbf{C}+\epsilon\boldsymbol{\Xi}^{u,v}$ on the original propagation matrix $\mathbf{C}$, the sensitivity of the Linear GCN model output $f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))$ as defined in Definition D.1 can be characterized by:
$\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})$ $:=~{}\frac{\mathrm{d}f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\mathrm{d}\epsilon}\Bigr{|}_{\epsilon=0}$
$=~{}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}\mathbf{H}^{-1}\Bigl{(}\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr{)},$
@@
The edge perturbation $\epsilon\boldsymbol{\Xi}$ influences $f_{\mathrm{LGN}}$ from two different aspects: propagation matrix and optimal model weights. Thus, the influence function can be derived by chain rule as follows:
$\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})$ $=~{}\frac{\mathrm{d}f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\mathrm{d}\epsilon}\Bigr{|}_{\epsilon=0}$
$=(\frac{\partial f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\partial(\mathbf{C}+\epsilon\boldsymbol{\Xi})}\cdot\frac{\partial(\mathbf{C}+\epsilon\boldsymbol{\Xi})}{\partial\epsilon})\Bigr{|}_{\epsilon=0}$
$\quad\quad\quad+(\frac{\partial f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\partial\mathbf{w}^{\star}(\epsilon)}\cdot\frac{\partial\mathbf{w}^{\star}(\epsilon)}{\partial\epsilon})\Bigr{|}_{\epsilon=0}$
$=~{}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}+(\frac{\partial f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\boldsymbol{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\partial\mathbf{w}^{\star}(\epsilon)}\cdot\frac{\partial\mathbf{w}^{\star}(\epsilon)}{\partial\epsilon})\Bigr{|}_{\epsilon=0}$
$=~{}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}(\frac{\partial\mathbf{w}^{\star}(\epsilon)}{\partial\epsilon})\Bigr{|}_{\epsilon=0}$
$=~{}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}\mathbf{H}^{-1}\Bigl{(}\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\boldsymbol{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\boldsymbol{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr{)}.$
@@
Next, we present two immediate corollaries of this theorem and determine the influence of a single edge on a single node’s output.
@@
Let $\mathbf{Z}:=\mathbf{C}\mathbf{X}$ and $\mathbf{z}_{l}$ is the $l$-th row of $\mathbf{Z}$. Considering the influence of one specific edge $(v_{i},v_{j})$ on undirected graphs (i.e., $\boldsymbol{\Xi}=\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\mathbf{e}_{j}\mathbf{e}_{i}^{\top}$), the influence function for the model output is:
@@
This follows from basic algebra and the fact that $\mathrm{vec}(\mathbf{A}\mathbf{X}\mathbf{B})=(\mathbf{B}\otimes\mathbf{A})\mathrm{vec}(\mathbf{X})$. ∎
Corollary D.11 (Single Edge Influence on Single Node’s Model Output, Linear GCN Case).
Let $\mathbf{Z}:=\mathbf{C}\mathbf{X}$ and $\mathbf{z}_{l}$ is the $l$-th row of $\mathbf{Z}$. Considering the influence of one specific edge $(v_{i},v_{j})$ on undirected graphs (i.e., $\boldsymbol{\Xi}=\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\mathbf{e}_{j}\mathbf{e}_{i}^{\top}$), the influence function for the model output for specific node $v_{k}$ is:
$\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})_{k}=$ $~{}\boldsymbol{1}\{v_{k}=v_{i}\}\cdot(\mathbf{x}^{\top}_{v}\mathbf{w}^{\star})+\boldsymbol{1}\{v_{k}=v_{j}\}\cdot(\mathbf{x}^{\top}_{u}\mathbf{w}^{\star})+\mathbf{q}^{\top}\mathbf{H}^{-1}\mathbf{z}_{k}$
$=$ $~{}\underbrace{\boldsymbol{1}{\{v_{k}=v_{i}\}}\cdot(\mathbf{x}^{\top}_{i}\mathbf{w}^{\star})+\boldsymbol{1}{\{v_{k}=v_{j}\}}\cdot(\mathbf{x}^{\top}_{j}\mathbf{w}^{\star})}_{\mathrm{edge~{}influence}}$
$~{}~{}~{}~{}-\underbrace{\langle(\mathbf{x}_{j}^{\top}\mathbf{w}^{\star})\mathbf{z}_{i}+(\mathbf{x}_{i}^{\top}\mathbf{w}^{\star})\mathbf{z}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{magnitude~{}weight~{}influence}}+\underbrace{\langle(y_{j}-\mathbf{z}_{j}^{\top}\mathbf{w}^{\star})\mathbf{x}_{i}+(y_{i}-\mathbf{z}_{i}^{\top}\mathbf{w}^{\star})\mathbf{x}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{error~{}weight~{}influence}},$
@@
This follows from basic algebra, the definition of $z$, and the definition of inner product w.r.t. a PSD matrix ($\langle\cdot,~{}\cdot\rangle_{\mathbf{A}}$). ∎
Appendix E Preliminary Experiments
@@
E.1 Probability Similarity Gap
Dataset Unlearn Method $\mathrm{ProbSim}(\mathcal{Q}^{-})$ $\mathrm{ProbSim}(\mathcal{Q}^{+}_{\mathrm{un}})$ $\mathrm{ProbSim}(\mathcal{Q}^{+}_{\mathrm{mem}})$
@@ E.1 Probability Similarity Gap
Table 2: Average black-box probability similarity between different types of edges. We consider three types of edges: negative edges ($\mathcal{Q}^{-}$), unlearned edges ($\mathcal{Q}^{+}_{\mathrm{un}}$), and other membership edges ($\mathcal{Q}^{+}_{\mathrm{mem}}$). The similarity measure is based on JS Divergence.
@@
$\mathrm{ProbSim}(\mathcal{Q}^{-})<\mathrm{ProbSim}(\mathcal{Q}^{+}_{\mathrm{un}})<\mathrm{ProbSim}(\mathcal{Q}^{+}_{\mathrm{mem}}).$
@@
E.2 Confidence Pitfall
(a) GIF
(b) CEU
(c) GA
Figure 5: The relation between average model confidence and distance to unlearned edges on Cora.
(a) GIF
(b) CEU
(c) GA
Figure 6: The relation between average model confidence and distance to unlearned edges on Citeseer.
@@ E.2 Confidence Pitfall
For all datasets, nodes directly connected to unlearned edges (distance = 0) exhibit substantially lower confidence than their immediate neighbors (distance $\geq$ 1). This aligns with our theoretical analysis in Section C, which shows that the influence of edge unlearning on the endpoint is significantly greater than on other nodes. This empirical finding suggests that the confidence gap between a node and its neighborhood is a strong indicator of its connection to an unlearned edge. Consequently, an attacker can adjust the similarity threshold accordingly to enhance membership inference accuracy.
@@ Appendix F Experimental Settings
All victim models use an embedding size of 16 and are trained for 100 epochs. For learning rates and weight decays, we follow the settings of GIF wu2023gif . All models are optimized using the Adam optimizer kingma2015adam . For all neural attack models, including StealLink he2021stealing , MIA-GNN olatunji2021membership , and TrendAttack, we follow the settings of StealLink he2021stealing , using a hidden dimension of 64 and a 2-layer MLP to encode features. All membership inference models are trained using learning rate = 0.01, weight decay = 0.0001, and the Adam optimizer kingma2015adam . They are trained with the binary cross-entropy loss.
Unlearning Method Settings.
We now supplement the missing details of our unlearning methods.
@@
  • GIF wu2023gif : The number of estimation iterations is 100 and the damping factor is 0. On the Cora and Citeseer datasets, the scale factor $\lambda$ is set to 500, following the official implementation. On the Pubmed and LastFM-Asia datasets, we search $\lambda$ in $\{10^{1},10^{2},10^{3},10^{4},\ldots\}$, following the original search space. This ensures that the unlearned models maintain utility and produce meaningful outputs. We select the smallest $\lambda$ that results in a meaningful model and find that $\lambda=10^{3}$ works well for Pubmed and $\lambda=10^{4}$ for LastFM-Asia.
@@
  • CEU wu2023certified : For a fair comparison, we use the same parameter search space as GIF for all influence-function-related parameters, including iterations, damping factor, and $\lambda$. For the noise variance, we search in $\{0.1,0.5,0.01,0.005,0.001,0.0005,0.0001\}$ to ensure meaningful unlearning performance.
@@
  • GA wu2024graphguard : We follow the official GraphGuard settings and use 1 gradient ascent epoch. We tune the unlearning magnitude $\alpha$ in $\{1,0.5,0.1,0.05,0.01,0.005,0.001\}$ and find that $\alpha=0.5$ gives the best results.
Baselines.
Details for our baselines are as follows:
@@
  • GraphSAGE hamilton2017inductive : We use a simple link prediction model trained on the shadow graph to perform link prediction on the unlearned graph. The embedding size is 64, with 2 layers. The negative-to-positive link ratio is 1:1.
@@
  • NCN wang2024neural : This is a state-of-the-art link prediction method. We follow the official settings from their code repository. We search across the provided backbones: GCN kipf2017semi , GIN xu2019powerful , and SAGE wang2023inductive . The embedding dimension is set to the default value of 256. The batch size is 16, the learning rate is 0.01, and the number of training epochs is 100.
@@
  • MIA-GNN olatunji2021membership : Since no official implementation is available, we follow the settings described in the paper. We re-implement the model using a 2-layer MLP feature extractor with a hidden dimension of 64. We use another 2-layer MLP with a dimension of 16 as the link predictor. Training uses binary cross-entropy loss and the same setup as StealLink, ensuring a fair comparison.
@@
  • StealLink he2021stealing : This is a representative membership inference attack. We use its strongest variant, Attack-7, which has access to the shadow dataset and partial features and connectivity from the target dataset. We follow the official codebase and use all 8 distance metrics (e.g., cosine, Euclidean). For the probability metric, we search among entropy, KL divergence, and JS divergence, and report the best result. The reference model is a 2-layer MLP with an embedding size of 16.
@@
  • GroupAttack zhang2023demystifying : This is a recent membership inference attack that relies solely on hard labels and thresholding. We search the threshold hyperparameter $\alpha$ on the shadow dataset in the range $[0,1]$ with a step size of 0.05.
Reproducibility.
@@
All experiments are conducted using Python 3.12.2, PyTorch 2.3.1+cu121, and PyTorch Geometric 2.5.3. The experiments run on a single NVIDIA RTX A6000 GPU. The server used has 64 CPUs, and the type is AMD EPYC 7282 16-Core CPU. All source code is provided in the supplementary materials.

### v2

@@ D.3. Edge Influence on Model Output
Lemma D.9 (Edge Influence on Model Output, Linear GCN Case).
Let $\bm{\Xi}^{\Delta\mathcal{E}}\in\mathbb{R}^{n\times n}$ be a perturbation matrix as defined in Definition D.6. Let $\epsilon\in\mathbb{R}$ be a perturbation magnitude and $\mathbf{w}^{\star}(\epsilon)$ be the optimal model weight after perturbation. Considering an edge perturbation $\mathbf{C}(\epsilon):=\mathbf{C}+\epsilon\bm{\Xi}^{u,v}$ on the original propagation matrix $\mathbf{C}$, the sensitivity of the Linear GCN model output $f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))$ as defined in Definition D.1 can be characterized by:
$~\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})$
$:=$ $~\frac{\mathrm{d}f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\mathrm{d}\epsilon}\Bigr|_{\epsilon=0}$
$=$ $~\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}\mathbf{H}^{-1}\Bigl(\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr),$
@@
The edge perturbation $\epsilon\bm{\Xi}$ influences $f_{\mathrm{LGN}}$ from two different aspects: propagation matrix and optimal model weights. Thus, the influence function can be derived by chain rule as follows:
$~\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})$
$=$ $~\frac{\mathrm{d}f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\mathrm{d}\epsilon}\Bigr|_{\epsilon=0}$
$=$ $~(\frac{\partial f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\partial(\mathbf{C}+\epsilon\bm{\Xi})}\cdot\frac{\partial(\mathbf{C}+\epsilon\bm{\Xi})}{\partial\epsilon})\Bigr|_{\epsilon=0}$
$\quad\quad\quad+(\frac{\partial f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\partial\mathbf{w}^{\star}(\epsilon)}\cdot\frac{\partial\mathbf{w}^{\star}(\epsilon)}{\partial\epsilon})\Bigr|_{\epsilon=0}$
$=$ $~\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}+(\frac{\partial f_{\mathrm{LGN}}(\mathbf{C}+\epsilon\bm{\Xi},\mathbf{X};\mathbf{w}^{\star}(\epsilon))}{\partial\mathbf{w}^{\star}(\epsilon)}\cdot\frac{\partial\mathbf{w}^{\star}(\epsilon)}{\partial\epsilon})\Bigr|_{\epsilon=0}$
$=$ $~\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}(\frac{\partial\mathbf{w}^{\star}(\epsilon)}{\partial\epsilon})\Bigr|_{\epsilon=0}$
$=$ $~\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}+\mathbf{C}\mathbf{X}\mathbf{H}^{-1}\Bigl(\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{y}-\mathbf{X}^{\top}\mathbf{C}^{\top}\bm{\Xi}\mathbf{X}\mathbf{w}^{\star}-\mathbf{X}^{\top}\bm{\Xi}^{\top}\mathbf{C}\mathbf{X}\mathbf{w}^{\star}\Bigr).$
@@
Next, we present an immediate corollary of this lemma and determine the influence of a single edge on all nodes’ output.
@@
Let $\mathbf{Z}:=\mathbf{C}\mathbf{X}$ and $\mathbf{z}_{l}$ is the ll‐th row of $\mathbf{Z}$. Considering the influence of one specific edge $(v_{i},v_{j})$ on undirected graphs (i.e., $\bm{\Xi}=\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\mathbf{e}_{j}\mathbf{e}_{i}^{\top}$), the influence function for the model output is:
@@
This follows from basic algebra and the fact that
$\mathrm{vec}(\mathbf{A}\mathbf{X}\mathbf{B})=(\mathbf{B}\otimes\mathbf{A})\mathrm{vec}(\mathbf{X})$. ∎
Then, taking an element $\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})_{k}$ from the vector-form influence function over all model outputs $\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})$, we obtain the final result on the influence of a single edge on a single node’s output.
Theorem D.11 (Single Edge Influence on Single Node’s Model Output, Linear GCN Case).
Let $\mathbf{Z}:=\mathbf{C}\mathbf{X}$ and $\mathbf{z}_{l}$ is the ll‐th row of $\mathbf{Z}$. Considering the influence of one specific edge $(v_{i},v_{j})$ on undirected graphs (i.e., $\bm{\Xi}=\mathbf{e}_{i}\mathbf{e}_{j}^{\top}+\mathbf{e}_{j}\mathbf{e}_{i}^{\top}$), the influence function for the model output for specific node $v_{k}$ is:
$\mathcal{I}_{\epsilon}(f_{\mathrm{LGN}})_{k}=$ $~\bm{1}\{v_{k}=v_{i}\}\cdot(\mathbf{x}^{\top}_{v}\mathbf{w}^{\star})+\bm{1}\{v_{k}=v_{j}\}\cdot(\mathbf{x}^{\top}_{u}\mathbf{w}^{\star})$
$\quad\quad+\mathbf{q}^{\top}\mathbf{H}^{-1}\mathbf{z}_{k}$
$=$ $~\underbrace{\bm{1}{\{v_{k}=v_{i}\}}\cdot(\mathbf{x}^{\top}_{i}\mathbf{w}^{\star})+\bm{1}{\{v_{k}=v_{j}\}}\cdot(\mathbf{x}^{\top}_{j}\mathbf{w}^{\star})}_{\mathrm{edge~influence}}$
$~~~~-\underbrace{\langle(\mathbf{x}_{j}^{\top}\mathbf{w}^{\star})\mathbf{z}_{i}+(\mathbf{x}_{i}^{\top}\mathbf{w}^{\star})\mathbf{z}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{magnitude~weight~influence}}$
$~~~~+\underbrace{\langle(y_{j}-\mathbf{z}_{j}^{\top}\mathbf{w}^{\star})\mathbf{x}_{i}+(y_{i}-\mathbf{z}_{i}^{\top}\mathbf{w}^{\star})\mathbf{x}_{j},\mathbf{z}_{k}\rangle_{\mathbf{H}^{-1}}}_{\mathrm{error~weight~influence}},$
@@
This follows from Corollary D.10, basic algebra, the definition of $z$, and the definition of inner product w.r.t. a PSD matrix ($\langle\cdot,~\cdot\rangle_{\mathbf{A}}$). ∎
Appendix E Preliminary Study
@@
E.1. Probability Similarity Gap
Table 2. Average black-box probability similarity between different types of edges. We consider three types of edges: negative edges ($\mathcal{Q}^{-}$), unlearned edges ($\mathcal{Q}^{+}_{\mathrm{un}}$), and other membership edges ($\mathcal{Q}^{+}_{\mathrm{mem}}$). The similarity measure is based on JS Divergence.
[…]
@@ E.1. Probability Similarity Gap
[…]
@@
E.2. Confidence Pitfall
(a) GIF results.
(b) CEU results.
(c) GA results.
Figure 6. The relation between average model confidence and distance to unlearned edges on Cora.
(a) GIF results.
(b) CEU results.
(c) GA results.
Figure 7. The relation between average model confidence and distance to unlearned edges on Citeseer.
@@ E.2. Confidence Pitfall
For all datasets, nodes directly connected to unlearned edges (distance = 0) exhibit substantially lower confidence than their immediate neighbors (distance $\geq$ 1). This aligns with our theoretical analysis in Section 5.1, which shows that the influence of edge unlearning on the endpoint is significantly greater than on other nodes. This empirical finding suggests that the confidence gap between a node and its neighborhood is a strong indicator of its connection to an unlearned edge. Consequently, an attacker can adjust the similarity threshold accordingly to enhance membership inference accuracy.
@@ Appendix F Experimental Settings
Model Parameters. All victim models use an embedding size of 16 and are trained for 100 epochs. For learning rates and weight decays, we follow the settings of GIF (wu2023gif). All models are optimized using the Adam optimizer (kingma2015adam). For all neural attack models, including StealLink (he2021stealing), MIA-GNN (olatunji2021membership), and TrendAttack, we follow the settings of StealLink (he2021stealing), using a hidden dimension of 64 and a 2-layer MLP to encode features. All membership inference models are trained using learning rate = 0.01, weight decay = 0.0001, and the Adam optimizer (kingma2015adam). They are trained with the binary cross-entropy loss.
Unlearning Method Settings. We now supplement the missing details of our unlearning methods.
@@
  • GIF (wu2023gif): The number of estimation iterations is 100 and the damping factor is 0. On the Cora and Citeseer datasets, the scale factor $\lambda$ is set to 500, following the official implementation. On the Pubmed and LastFM-Asia datasets, we search $\lambda$ in $\{10^{1},10^{2},10^{3},10^{4},\ldots\}$, following the original search space. This ensures that the unlearned models maintain utility and produce meaningful outputs. We select the smallest $\lambda$ that results in a meaningful model and find that $\lambda=10^{3}$ works well for Pubmed and $\lambda=10^{4}$ for LastFM-Asia.
@@
  • CEU (wu2023certified): For a fair comparison, we use the same parameter search space as GIF for all influence-function-related parameters, including iterations, damping factor, and $\lambda$. For the noise variance, we search in $\{0.1,0.5,0.01,0.005,0.001,0.0005,0.0001\}$ to ensure meaningful unlearning performance.
@@
  • GA (wu2024graphguard): We follow the official GraphGuard settings and use 1 gradient ascent epoch. We tune the unlearning magnitude $\alpha$ in $\{1,0.5,0.1,0.05,0.01,0.005,0.001\}$ and find that $\alpha=0.5$ gives the best results.
Baselines. Details for our baselines are as follows:
@@
  • GraphSAGE (hamilton2017inductive): We use a simple link prediction model trained on the shadow graph to perform link prediction on the unlearned graph. The embedding size is 64, with 2 layers. The negative-to-positive link ratio is 1:1.
@@
  • NCN (wang2024neural): This is a state-of-the-art link prediction method. We follow the official settings from their code repository. We search across the provided backbones: GCN (kipf2017semi), GIN (xu2019powerful), and SAGE (wang2023inductive). The embedding dimension is set to the default value of 256. The batch size is 16, the learning rate is 0.01, and the number of training epochs is 100.
@@
  • MIA-GNN (olatunji2021membership): Since no official implementation is available, we follow the settings described in the paper. We re-implement the model using a 2-layer MLP feature extractor with a hidden dimension of 64. We use another 2-layer MLP with a dimension of 16 as the link predictor. Training uses binary cross-entropy loss and the same setup as StealLink, ensuring a fair comparison.
@@
  • StealLink (he2021stealing): This is a representative membership inference attack. We use its strongest variant, Attack-7, which has access to the shadow dataset and partial features and connectivity from the target dataset. We follow the official codebase and use all 8 distance metrics (e.g., cosine, Euclidean). For the probability metric, we search among entropy, KL divergence, and JS divergence, and report the best result. The reference model is a 2-layer MLP with an embedding size of 16.
@@
  • GroupAttack (zhang2023demystifying): This is a recent membership inference attack that relies solely on hard labels and thresholding. We search the threshold hyperparameter $\alpha$ on the shadow dataset in the range $[0,1]$ with a step size of 0.05.
@@
Reproducibility. All experiments are conducted using Python 3.12.2, PyTorch 2.3.1+cu121, and PyTorch Geometric 2.5.3. The experiments run on a single NVIDIA RTX A6000 GPU. The server used has 64 CPUs, and the type is AMD EPYC 7282 16-Core CPU. All source code is provided in the supplementary materials.